OpenStudy (anonymous):

Find the sum of:

OpenStudy (anonymous):

\[2\sqrt{6} and \sqrt{24}\]

OpenStudy (anonymous):

sqrt(24) can be divided into sqrt(4) and sqrt(6). The sqrt of 4 is 2. so you have 2\[\sqrt{6}+2\sqrt{6}\]

OpenStudy (anonymous):

which equals... :)

OpenStudy (anonymous):

Uhm.. 2 *square root* 6 + 4 *square root* 3??

OpenStudy (anonymous):

sorry, my stuff got kinda out of order... \[2\sqrt{6}+2\sqrt{6}\] This equals\[4\sqrt{6}\]

OpenStudy (anonymous):

Lol. Okay. Thanks for the help brycoope(: